OpenStudy (anonymous):

If the scale on a map is 1 inch = 850 miles, how many inches would represent 6460 miles.

OpenStudy (cj49):

6460/850= ? inches

OpenStudy (anwaarullah):

1 inche : 850 x inches ; 6460 cross mutipication x*850= 6460*1 x=6460/850 x=7.6